Effect of proposals on divisional performance

A condensed income statement for the Commercial Division of Maxell Manufacturing Inc. for the year ended December 31, 20Y9, is as follows:

Sales $3,500,000
Cost of goods sold (2,480,000)
Gross profit $1,020,000
Operating expenses (600,000)
Operating income $420,000
Invested assets $2,500,000

Assume that the Commercial Division received no allocations from support departments.

The president of Maxell Manufacturing has indicated that the division’s return on a $2,500,000 investment must be increased to at least 21% by the end of the next year if operations are to continue. The division manager is considering the following three proposals:

Proposal 1: Transfer equipment with a book value of $312,500 to other divisions at no gain or loss and lease similar equipment. The annual lease payments would exceed the amount of depreciation expense on the old equipment by $105,000. This increase in expense would be included as part of the cost of goods sold. Sales would remain unchanged.

Proposal 2: Purchase new and more efficient machining equipment and thereby reduce the cost of goods sold by $560,000 after considering the effects of depreciation expense on the new equipment. Sales would remain unchanged, and the old equipment, which has no remaining book value, would be scrapped at no gain or loss. The new equipment would increase invested assets by an additional $1,875,000 for the year.

Proposal 3: Reduce invested assets by discontinuing a product line. This action would eliminate sales of $595,000, reduce cost of goods sold by $406,700, and reduce operating expenses by $175,000. Assets of $1,338,000 would be transferred to other divisions at no gain or loss.

Required:

1. Using the DuPont formula for return on investment, determine the profit margin, investment turnover, and return on investment for the Commercial Division for the past year. If required, round your answers to two decimal place.

Commercial Division
Profit margin %
Investment turnover
ROI %

2. Prepare condensed estimated income statements and compute the invested assets for each proposal.

Maxell Manufacturing Inc.—Commercial Division
Estimated Income Statements
For the Year Ended December 31, 20Y9
Proposal 1 Proposal 2 Proposal 3
Sales $ $ $
Cost of goods sold
Gross profit $ $ $
Operating expenses
Operating income $ $ $
Invested assets $ $ $

3. Using the DuPont formula for return on investment, determine the profit margin, investment turnover, and return on investment for each proposal. Round the investment turnover and return on investment to one decimal place.

Proposal Profit margin Investment turnover ROI
Proposal 1 % %
Proposal 2 % %
Proposal 3 % %

4. Which of the three proposals would meet the required 21% return on investment?

Proposal 1
Proposal 2
Proposal 3

5. If the Commercial Division were in an industry where the profit margin could not be increased, how much would the investment turnover have to increase to meet the president’s required 21% return on investment? Enter your increase in investment turnover answer as a percentage of current investment turnover. Round interim calculations (including previously calculated) and final answer to one decimal place.
%
Answer #1

Solution 1) Maxell manufacturing Inc. Return on Investment = profit Margen & Investment Turnover Commercial Division i Roi =2480000 + 105000 = 2585000 2500000 - 312500 = 218.750D 2480 000 - 560000 = 1920000 2500000 + 1875000 = 4375000 3500000 - 5954) proposal 3 would yield a rate of 35 %, proposal 2 would field a return rate of on Investment of return on Investment of 27- 5 Rol = profit Margen y Investment Turnover . 21% = 12% * Investment Turnover - Required Investment Turnover = 21% 12% = 1.